OpenStudy (anonymous):

When two resistors are connected in parallel, the effective resistance is given by the expression 1/(1/R1+1/R2) Simplify the complex fraction

OpenStudy (ipwnbunnies):

\[\frac{1}{\frac{1}{R_{1}} + \frac{1}{R_{2}}}\] Hint: Rewrite the denominator of this complex fraction. You want to find a common denominator for the denominator of the fraction.